# Examining Post-Training Quantization for Mixture-of-Experts: A Benchmark

### v1

@@
Abstract
Large Language Models (LLMs) have become foundational in the realm of natural language processing, demonstrating performance improvements as model sizes increase. The Mixture-of-Experts (MoE) approach offers a promising way to scale LLMs more efficiently by using fewer computational FLOPs through sparse activation. However, it suffers from significant memory overheads, necessitating model compression techniques. Post-training quantization, a popular method for model compression, proves less effective when directly applied to MoE models due to MoE’s overlooked inherent sparsity. This paper explores several MoE structure-aware quantization heuristics, ranging from coarse to fine granularity, from MoE block to individual linear weight. Our investigations reveal critical principles: different MoE structures (i.e., blocks, experts, linear layers) require varying numbers of weight bits for effective and efficient quantization. Conclusions are supported by extensive benchmarking across two representative MoE models and six tasks. We further introduce novel enhancements to more accurately identify the most critical weights in MoE quantization that necessitate higher bit allocations, including the linear weight outlier scorer and MoE block scorer. Additionally, subsequent experiments validate our findings in the context of both weight and activation quantization. Our code for reproducing all our experiments is provided at https://github.com/UNITES-Lab/moe-quantization.
1 Introduction
@@
Large Language Models (LLMs) have achieved remarkable success in various natural language processing tasks, such as language understanding, reasoning, and generation, demonstrating superior performance and adaptability [4, 13, 15, 23, 29]. However, the rapid growth in model size, with state-of-the-art LLMs containing billions of parameters, poses significant challenges to computational resources and memory consumption [2, 18, 28]. The Mixture of Experts (MoE) [27] architecture has emerged as a promising solution to address these challenges. MoE allows for the scaling up of LLMs while maintaining roughly constant FLOPs. By incorporating multiple expert networks and employing a sparse gating mechanism, MoE achieves efficient computation, enabling the development of larger models within the constraints of limited computational resources [6, 7, 14].
@@
Despite its advantages, MoE suffers from extensive memory costs, which hinder its practical deployment and widespread adoption. For example, the Mixtral-8x7B [14] MoE model takes around $180$ GB memory while only $28$ GB parameters are activated for each input token [^1] This is evaluated in full precision (float32323232).. Model compression techniques tailored to MoE architectures are essential to address this issue. Existing MoE compression methods can be categorized into two main approaches: merging and pruning. Expert merging, such as MC-MoE[17], aims to reduce the memory footprint by combining similar experts based on routing policy and compressing the resulting model using low-rank decomposition. On the other hand, expert pruning, such as task-specific pruning [5], focuses on identifying and removing the least important experts or connections based on their contribution to a specific task. However, these approaches ① necessitate model retraining, which is both extremely costly and time-consuming, particularly for state-of-the-art MoE LLMs of billion-size scale, and ② operate under task-specific settings, which limits their practicality for real-world applications.
@@
Post-training quantization has emerged as a promising compression method widely applied to dense LLM models. Recent works, such as GPTQ [8], which adapts quantization intervals based on the Hessian information, SmoothQuant [18], which jointly quantizes the model weight and activation by offline migrating the activation outliers, have demonstrated the effectiveness of post-training quantization for LLMs toward $4$ bits compression.
However, directly applying existing quantization methods to MoE models in a more extreme quantization setting, e.g. under $3$ bits, leads to suboptimal results, potentially due to the overlooked sparsity nature of the MoE architecture. The sparse activation patterns and the dynamic routing mechanism in MoE pose unique challenges and opportunities for quantization, requiring novel approaches to utilize it effectively. The sparse expert activations in MoE models exhibit different statistical properties methodologies compared to dense activations, making conventional quantization methods difficult. Moreover, the dynamic routing mechanism, which selects a subset of experts for each input token, introduces additional complexity in terms of quantizing the routing weights and maintaining the sparsity pattern during inference. This yields the primary question to be explored:
(Q) Can we leverage the sparsity nature of MoE architecture to establish more efficient and effective coarse-grained mixed-precision MoE quantization methods?
To answer (Q), we explore a wide range of MoE structure-aware quantization heuristics, ranging from coarse to fine granularity. We conduct a detailed comparative analysis of each of them, revealing critical principles: different MoE structures (i.e., blocks, experts, linear layers) require varying numbers of weight bits for effective and efficient quantization. Extended from the gained insights, we propose methods to further improve the efficiency and effectiveness of mixed-precision quantization, including linear weight quantization scorer and MoE block quantization scorer.
In summary, our key contributions are listed below:
  1. 1.
    We establish the first benchmark for post-training quantization specifically designed for the Mixture-of-Experts architecture. This benchmark encompasses investigations into four critical MoE-related heuristics, evaluations across two MoE LLMs, six benchmark tasks, and a combination of both weight and activation quantization.
  2. 2.
    Our benchmark study uncovers a range of previously unexplored quantization principles and insights for MoE. These insights include empirical rules supporting optimal bit allocation strategies, highlighting the trade-offs such us those between attention and FFNN layers, and among different experts.
  3. 3.
    Leveraging the insights from our benchmark study, we introduce novel enhancements to improve existing heuristics. These include the development of linear-weight and MoE block scorers to identify the most critical components of the MoE model, thereby guiding more effective quantization bit assignments.
2 Related Works
Mixture-of-Experts.
@@ 2 Related Works
The Mixture-of-Experts (MoE) approach [27] enhances neural network scalability by using router networks to activate model segments according to input tokens selectively. As the dominant architecture in NLP, numerous efforts have adapted feed-forward neural networks (FFNNs) within Transformers to incorporate MoE layers, constructing MoE language models[6, 7, 14]. Additionally, several variants of the standard MoE architecture exist. For example, DeepSeek-MoE [6] employs numerous finely segmented experts and designates a select few as shared experts to capture common knowledge. MoE’s application in LLMs is widely acknowledged for its superior generative abilities and remarkable computing efficiency [3, 6, 7, 14, 16, 25]. The recent work Mixtral [14] illustrates that MoE can match the performance of equivalent full-parameter LLMs while utilizing far fewer active parameters. However, MoE suffers from significant memory overhead issues, posing challenges to its efficient deployment [17].
MoE Compression.
@@
MoE models benefit from reduced FLOPs but are constrained by their significant memory overhead. Current works to reduce the memory overhead of MoE models mainly focus on reducing the number of experts. An earlier approach [5] involves pruning non-essential experts for a specific downstream task during fine-tuning, utilizing statistics based on cumulative usage frequency. Another method, MC-SMoE [17], introduces a pipeline that identifies and groups similar experts, subsequently merging them and further decomposing the merged expert into low-rank components within each group. However, these approaches are developed under task-specific fine-tuning settings and do not explore the development of the MoE compression towards a general post-training model.
Post-Training Quantization.
@@
Post-training quantization reduces computational and storage demands by converting pre-trained models from high-precision to lower-precision formats without extensive retraining [9, 8]. It has been widely applied to LLMs, optimizing them for deployment on resource-constrained devices. Techniques like layer-wise quantization and mixed-precision schemes are designed for minimal performance degradation while reducing model size and computational requirements efficiently [19, 24, 26]. Recent methods such as SmoothQuant [30], GPTQ [8], AWQ [18], and address specific challenges for LLMs. SmoothQuant [30] ensures smooth precision transitions across layers, reducing quantization errors and maintaining performance. GPTQ [8] employs layer-wise and mixed-precision quantization to balance efficiency and accuracy. AWQ [18] adapts to weight sensitivity, preserving critical weights’ precision while aggressively quantizing less sensitive ones. These advancements in PTQ enable significant reductions in computational and storage requirements while preserving LLM performance.
3 Reviewing Quantization and MoE
3.1 Quantization Method
@@
The primary objective of this work is to benchmark several MoE-related heuristics combined with established LLM quantization techniques. Given that the substantial memory overhead of MoE models predominantly originates from their weights, we adopt GPTQ [8], a popular weight quantization method. GPTQ executes layer-by-layer weight quantization by addressing a specific reconstruction problem for each layer. Specifically, let $\mathbf{W}$ represent the weights of a linear layer and $\mathbf{X}$ denote the input to that layer derived from a small subset of calibration data, the reconstruction problem is defined as follows:
$\texttt{argmin}_{\mathbf{\widehat{W}}},||\mathbf{W}\mathbf{X}-\mathbf{\widehat{W}}\mathbf{X}||_{2}^{2}.$ (1)
@@ 3.2 Mixture-of-Experts
$\mathtt{FFNN}_{\text{MoE}}^{l}(\mathbf{X})=\sum_{i=1}^{l}\mathcal{G}(\mathbf{W}_{l}\mathbf{X})\cdot\mathtt{FFNN}_{i}^{l}(\mathbf{X}),$ (2)
@@
3.3 Expert Usage as A Heuristic
As the routing of experts in MoE models is not ideally balanced, expert usage frequency and its variants have emerged as prevalent heuristics for measuring the importance of different experts within an MoE block [5, 17]. For instance, task-specific expert pruning proposed by [5] uses a criterion based on cumulatively calculated expert routing probabilities for pruning during fine-tuning on a specific task. In this paper, focusing on post-training quantization, we utilize the routing distribution from the calibration data as the heuristic for expert usage. Specifically, for the $l$-th MoE block, equipped with a routing matrix $\mathbf{W}l\in\mathbb{R}^{e\times d}$ and input hidden states $\mathbf{X}\in\mathbb{R}^{b\times d}$ from the calibration data, the expert usage heuristic is calculated as follows:
$\texttt{usage}=\texttt{normalize}\left(\sum_{i}\mathcal{G}(\mathbf{W}_{l}\mathbf{X}_{i})\right),$ (3)
@@ 3.3 Expert Usage as A Heuristic
Figure 1: Visualization of expert usage of the two MoE models used in this work. It is evaluated on the quantization calibration data, i.e., $512$ random $4096$ token sequences from the WikiText dataset [21].
4 Benchmark Post-Quantization Methods for MoE
@@ 4.1 Benchmark Setups
We select two representative MoE models for our benchmark evaluation, i.e., Mixtral-8x7B [14] and DeepSeek-MoE-16B-base [6]. Mixtral-8x7B substitutes every FFNN with a MoE block and has $8$ experts per MoE block with top-$2$ routing, while DeepSeed-MoE-16B-base uses a fine-grained MoE architecture by including $64$ experts with top-$6$ routing and $2$ shared experts per MoE block. Notably, the DeepSeek-MoE-16B-base model incorporates a dense architecture in its first transformer block while employing an MoE architecture in subsequent blocks for better training stability.
Quantization.
@@
We mainly focus on weight-only grouped mixed-precision quantization, though we also extend our experiments and conclusions to its combination with activation quantization in Section 5. The weight-only experiments utilize GPTQ [8], while those that combine weight and activation quantization utilize SmoothQuant [30], without loss of generality. Throughout this work, we use a group size of $128$. Our experiments emphasize an extreme quantization scenario, where most weights are quantized to either $2$ or $4$ bits.
Calibration and Evaluation Details.
@@
We use the calibration data consisting of $512$ random $4096$ token sequences from the WikiText dataset [21], following GPTQ [8]. Unlike previous literature that focuses on language modeling benchmarks [30, 18, 8], we evaluate all the methods on six popular LLM tasks for a practical benchmarking: WinoGrande [1], COPA [10], OpenBookQA (OBQA) [22], HellaSwag [31], and MMLU [11]. We report the performance on MMLU with $5$-shot and all others with zero-shot. All experiments are conducted with PyTorch on $3$ NVIDIA H100, and we utilize lm-evaluation-harness  [^2] https://github.com/EleutherAI/lm-evaluation-harness for the evaluation of all tasks.
4.2 Benchmark Results
We first evaluate several MoE heuristics quantization methods based on GPTQ on Mixtral-8x7B and DeepSeek-MoE-16B. We present our benchmark conclusions by answering the following research questions.

### v2

# QuantMoE-Bench: Examining Post-Training Quantization for Mixture-of-Experts

@@
Abstract.
Mixture-of-Experts (MoE) is a promising way to scale up the learning capacity of large language models. It increases the number of parameters while keeping FLOPs nearly constant during inference through sparse activation. Yet, it still suffers from significant memory overheads due to the vast parameter size, necessitating model compression techniques. Post-training quantization offers a powerful approach for model compression. Existing methods adopt a fixed quantization precision for the entire MoE model. This rigid setup can lead to suboptimal performance, without considering the inherent sparse structure. For example, MoE’s sparse routing mechanism leads to different activation patterns, where shared experts are accessed by all tokens while token-conditioned experts are selectively activated. This activation disparity suggests different quantization requirements, with consistently activated shared experts potentially needing higher precision to maintain model quality. In this paper, we study a fine-grained precision setup for MoE quantization. We explore MoE structure-aware quantization heuristics, ranging from coarse (e.g., MoE layers) to fine granularity (e.g., linear layers). Our investigations reveal critical principles, where different MoE structures require varying numbers of bits for effective quantization. Conclusions are supported by extensive benchmarking across two representative MoE models and six tasks including commonsense reasoning and natural language understanding. We further show that an MoE quantized in a fined-grained mixed precision achieved state-of-the-art $65.35\%$ performance on average compared to the baseline $64.30\%$ (i.e., GPTQ). Moreover, based on the findings, we introduce novel data-driven techniques for optimizing bit allocation in MoE quantization, including the outlier-aware linear layer scorer and MoE block importance predictor. Our code for reproducing all our experiments is provided at https://github.com/UNITES-Lab/moe-quantization.
Quantization, Mixture-of-Experts, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Neural networks
Figure 1. After our post-training quantization, the pre-trained MoE (a) is quantized into (b) with a mixed-precision based on structures, demonstrated with Mixtral-8x7B.
1. Introduction
@@
Large Language Models (LLMs) have achieved remarkable success in various natural language processing tasks, such as language understanding, reasoning, and generation, demonstrating superior performance and adaptability (Brown et al., 2020; Jiang et al., 2023; Kaplan et al., 2020; OpenAI et al., 2024; Touvron et al., 2023). However, the rapid growth in model size, with state-of-the-art LLMs containing billions of parameters, poses significant challenges to computational resources and memory consumption (Aminabadi et al., 2022; Lin et al., 2024; Shoeybi et al., 2020). The Mixture of Experts (MoE) (Shazeer et al., 2017) architecture has emerged as a promising solution to address the computation overhead. MoE allows for the scaling up of LLMs while maintaining roughly constant FLOPs. By incorporating multiple experts and employing a sparse gating mechanism, MoE achieves efficient computation, enabling the development of larger models within the constraints of limited computation (Dai et al., 2024).
@@
Despite its advantages, MoE still suffers from extensive memory costs due to its vast parameter size, which hinders its practical deployment and widespread adoption. For example, pretrained in 8-bit precision, the DeepSeek-V3 (Liu et al., 2024) MoE model takes around $1.3$ TB memory while only $74$ GB parameters are activated for each input token. Model compression techniques tailored to MoE architectures are essential to address this issue. Existing MoE compression methods can be categorized into two main approaches: merging and pruning. Expert merging, such as MC-MoE(Li et al., 2024), aims to reduce the memory footprint by combining similar experts based on routing policy and compressing the resulting model using low-rank decomposition. On the other hand, expert pruning, such as task-specific pruning (Chen et al., 2022), focuses on identifying and removing the least important experts or connections based on their contribution to a specific task. However, these approaches not only necessitate prohibitively expensive model retraining but also operate under task-specific settings, limiting their practicality across real-world applications.
@@
Post-training quantization is a family of model compression techniques that converts pre-trained model weights from high-precision formats (e.g., FP32323232) to lower-precision representations without model retraining or task-specific tuning. Recent works, such as GPTQ (Frantar et al., 2023a), which adapts quantization intervals based on the Hessian information, and SmoothQuant (Lin et al., 2024), which jointly quantizes the model weight and activation by offline migrating the activation outliers, have demonstrated the effectiveness of post-training quantization for dense LLMs toward $4$ bits compression. This advantage is particularly desired for MoE-based LLMs given their high training cost due to vast parameter sizes and prevailing deployment for various tasks (Jiang et al., 2024; Liu et al., 2024).
However, our experiments show that directly deploy those techniques for MoE models leads to subpar performance. Existing methods employ a uniform bit-width across all components of the MoE model. This one-size-fits-all approach fails to account for the inherent sparse structure of the MoE architecture. For example, the sparse expert activations in MoE models exhibit distinct statistical properties from dense activations, suggesting adaptive bit allocation among experts’ quantization. This yields the primary research question to be explored:
(RQ) Do different MoE structures require varying numbers of bits for effective quantization?
Our investigations reveal that different components in MoE require varying bit allocations, as shown in Figure 1. For example, shared experts and the first few MoE layers demand higher precision for effective quantization. Moreover, these findings naturally motivate two key questions: ($1$) How to identify the layers that are more sensitive to quantization; ($2$) How to systematically determine the importance of each MoE layer for bit allocation. To address these questions, we introduce novel data-driven techniques for optimizing bit allocation in MoE quantization, including the outlier-aware linear layer scorer that captures weight magnitude variations, and the MoE block importance predictor that leverages block-level activations patterns. Our key contributions are listed:
  1. (1)
    We establish the first benchmark for Mixture-of-Experts post-training quantization, i.e., QuantMoE-Bench. This benchmark encompasses investigations into four critical MoE-related heuristics by evaluating different quantization methods including GPTQ and SmoothQuant, and analyzes multiple bit allocation strategies on attention layers, FFNN layers, experts, and MoE blocks. Our evaluation covers two representative MoE LLMs and six benchmark tasks.
  2. (2)
    Our benchmark study uncovers critical MoE quantization principles: attention layers require higher precision than FFNNs, shared experts need more bits than token-conditioned experts ($4$-bit v.s. $2$-bit), and earlier MoE layers demand higher precision compared to later ones. These insights enable optimal bit allocation under constrained memory budgets while maintaining model performance.
  3. (3)
    Through extensive experiments, we demonstrate that our fine-grained mixed precision quantization approach achieves state-of-the-art performance, improving average task performance by $1.05\%$ compared to existing methods (i.e., GPTQ).
  4. (4)
    Leveraging the insights from our benchmark study, we introduce novel data-driven techniques to optimize bit allocation in MoE quantization. These include the development of outlier-aware linear layer scorer and MoE block importance predictor, which significantly improve the effectiveness of mixed-precision quantization by $0.97\%$.
2. Related Works
Figure 2. Visualization of expert usage of the two MoE models used in this work. It is profiled on our quantization calibration data, i.e., $512$ random $4096$ token sequences from the WikiText dataset (Merity et al., 2016).
@@ 2. Related Works
Mixture-of-Experts. The Mixture-of-Experts (MoE) approach (Shazeer et al., 2017) enhances neural network scalability by using router networks to activate model segments according to input tokens selectively. As the dominant architecture in NLP, numerous efforts have adapted feed-forward neural networks (FFNNs) within Transformers to incorporate MoE layers, constructing MoE language models(Dai et al., 2024; Fedus et al., 2022; Jiang et al., 2024). Additionally, several variants of the standard MoE architecture exist. For example, DeepSeek-MoE (Dai et al., 2024) employs numerous finely segmented experts and designates a select few as shared experts to capture common knowledge. MoE’s application in LLMs is widely acknowledged for its superior generative abilities and remarkable computing efficiency (Artetxe et al., 2022; Dai et al., 2024; Fedus et al., 2022; Jiang et al., 2024; Krajewski et al., 2024; Rajbhandari et al., 2022). The recent work Mixtral (Jiang et al., 2024) illustrates that MoE can match the performance of equivalent full-parameter LLMs while utilizing far fewer active parameters. However, MoE suffers from significant memory overhead issues, posing challenges to its efficient deployment (Li et al., 2024; Liu et al., 2024; Xue et al., 2024; Krajewski et al., 2024; Luo et al., 2024).
@@
MoE Compression. MoE models benefit from reduced FLOPs but are constrained by their significant memory overhead. Current works to reduce the memory overhead of MoE models mainly focus on reducing the number of experts. An earlier approach (Chen et al., 2022) involves pruning non-essential experts for a specific downstream task during fine-tuning, utilizing statistics based on cumulative usage frequency. Another method, MC-SMoE (Li et al., 2024), introduces a pipeline that identifies and groups similar experts, subsequently merging them and further decomposing the merged expert into low-rank components within each group. However, these approaches are developed under task-specific fine-tuning settings and do not explore the development of the MoE compression towards a general large langauge model.
@@
Post-Training Quantization. Post-training quantization reduces computational and storage demands by converting pre-trained models from high-precision to lower-precision formats without extensive retraining (Frantar et al., 2023b, a). It has been widely applied to LLMs, optimizing them for deployment on resource-constrained devices. Techniques like layer-wise quantization and mixed-precision schemes are designed for minimal performance degradation while reducing model size and computational requirements efficiently (Liu et al., 2023; Pan et al., 2023; Sharify et al., 2024). Recent methods such as SmoothQuant (Xiao et al., 2024), GPTQ (Frantar et al., 2023a), AWQ (Lin et al., 2024), and address specific challenges for LLMs. SmoothQuant (Xiao et al., 2024) ensures smooth precision transitions across layers, reducing quantization errors and maintaining performance. GPTQ (Frantar et al., 2023a) employs layer-wise and mixed-precision quantization to balance efficiency and accuracy. AWQ (Lin et al., 2024) adapts to weight sensitivity, preserving critical weights’ precision while aggressively quantizing less sensitive ones. These advancements in PTQ enable significant reductions in computing and storage while preserving LLM performance. However, their effectiveness on sparse MoE models is under-explored.
3. Preliminary
3.1. Quantization Method
@@
The primary objective of this work is to benchmark several MoE-related heuristics combined with established LLM quantization techniques. Given that the substantial memory overhead of MoE models predominantly originates from their weights, we adopt GPTQ (Frantar et al., 2023a), a popular weight quantization method. GPTQ executes layer-by-layer weight quantization by addressing a specific reconstruction problem for each layer. Specifically, let $\mathbf{W}$ be the weights of a linear layer and $\mathbf{X}$ be the input to that layer derived from a small subset of calibration data, the reconstruction problem is defined as:
(1) $\texttt{argmin}_{\mathbf{\widehat{W}}},||\mathbf{W}\mathbf{X}-\mathbf{\widehat{W}}\mathbf{X}||_{2}^{2}.$
@@ 3.2. Mixture-of-Experts
(2) $\mathtt{FFNN}_{\text{MoE}}^{l}(\mathbf{X})=\sum_{i=1}^{l}\mathcal{G}(\mathbf{W}_{l}\mathbf{X})\cdot\mathtt{FFNN}_{i}^{l}(\mathbf{X}),$
@@
3.3. Expert Usage as A Heuristic
As the routing of experts in MoE models is not ideally balanced, expert usage frequency and its variants have emerged as prevalent heuristics for measuring the importance of different experts within an MoE block (Chen et al., 2022; Li et al., 2024). For instance, task-specific expert pruning proposed by (Chen et al., 2022) uses a criterion based on cumulatively calculated expert routing probabilities for pruning during fine-tuning on a specific task. In this paper, focusing on post-training quantization, we utilize the routing distribution from the calibration data as the heuristic for expert usage. Specifically, for the $l$-th MoE block, equipped with a routing matrix $\mathbf{W}_{l}\in\mathbb{R}^{e\times d}$ and input hidden states $\mathbf{X}\in\mathbb{R}^{b\times d}$ from the calibration data, the expert usage heuristic is:
(3) $\texttt{usage}=\texttt{normalize}\left(\sum_{i}\mathcal{G}(\mathbf{W}_{l}\mathbf{X}_{i})\right),$
@@ 3.3. Expert Usage as A Heuristic
4. Benchmark Post-Quantizations for MoE
@@ 4.1. Benchmark Setups
MoE Models. We select two representative MoE models for our benchmark evaluation, i.e., Mixtral-8x7B (Jiang et al., 2024) and DeepSeek-MoE-16B-base (Dai et al., 2024). Mixtral-8x7B substitutes every FFNN with a MoE block and has $8$ experts per MoE block with top-$2$ routing, while DeepSeek-MoE-16B-base uses a fine-grained MoE architecture by including $64$ experts with top-$6$ routing and $2$ shared experts per MoE block. Notably, the DeepSeek-MoE-16B-base model applies a dense FFNN in its first transformer layer while employing an MoE architecture in subsequent blocks for better training stability.
@@
Quantization. We mainly focus on weight-only grouped mixed-precision quantization, though we also extend our experiments and conclusions to its combination with activation quantization in Section 5. We use GPTQ (Frantar et al., 2023a) for quantization, without loss of generality. Throughout this work, we use a group size of $128$. Our experiments emphasize an extreme quantization scenario, where most weights are quantized to either $2$ or $4$ bits.
@@
Calibration and Evaluation Details. We use the calibration data consisting of $512$ random $4096$ token sequences from the WikiText dataset (Merity et al., 2016), following GPTQ (Frantar et al., 2023a). Unlike previous literature that focuses on language modeling benchmarks (Xiao et al., 2024; Lin et al., 2024; Frantar et al., 2023a), we evaluate all the methods on six popular LLM tasks for a practical benchmarking: WinoGrande (ai2, 2019), COPA (Gordon et al., 2012), OpenBookQA (OBQA) (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019), and MMLU (Hendrycks et al., 2021). We report the performance on MMLU with $5$-shot and all others with zero-shot. All experiments are conducted with PyTorch on $8$ NVIDIA H100100100100s, and we utilize lm-evaluation-harness  [^1] https://github.com/EleutherAI/lm-evaluation-harness for the evaluation of all tasks.
4.2. Problem Formulation: Pareto-Optimal Bit Allocation
In this section, we formalize the intrinsic tradeoff between model performance and bit allocation in MoE quantization through the lens of Pareto optimization.
Given an MoE model with $n$ weight components (attention layers, MoE FFNN experts, etc.), we aim to find the optimal bit allocation strategy that balances model performance and memory efficiency. Let $\mathbf{b}=(b_{1},b_{2},\ldots,b_{n})$ represent a bit allocation vector, where $b_{i}\in\{2,4\}$ is the quantization precision for the $i$-th component. We formulate this as a bi-objective optimization problem:
(4) $\max_{\mathbf{b}}$ $\mathcal{P}(\mathbf{b})$
$\vspace{-10pt}\min_{\mathbf{b}}$ $\mathcal{B}(\mathbf{b})=\frac{1}{n}\sum_{i=1}^{n}b_{i}$
subject to $b_{i}\in\{2,4\},\quad\forall i\in\{1,2,\ldots,n\},$
where $\mathcal{P}(\mathbf{b})$ is the model performance and $\mathcal{B}(\mathbf{b})$ is the average bit-width across all components. In practice, we can reformulate this as a constrained optimization problem:
(5) $\max_{\mathbf{b}}$ $\mathcal{P}(\mathbf{b})$
subject to $\mathcal{M}(\mathbf{b})\leq M_{\text{budget}}$
$b_{i}\in\mathcal{B},\quad\forall i\in\{1,2,\ldots,N\},$
where $M_{\text{budget}}$ is the maximum allowed memory budget. This formulation allows us to find the bit allocation that maximizes performance while satisfying memory constraints.
Given the combinatorial nature of this problem and the difficulty in obtaining a closed-form expression for $\mathcal{P}(\mathbf{b})$, we approach it using the structure-aware heuristics described in the following sections, leveraging our observations about the varying importance of different MoE components. As shown in Figure 3, our mixed-precision approach effectively navigates the Pareto frontier, offering superior performance compared to uniform bit allocation.
Figure 3. Comparison of allocating more bits (i.e. $4$ bits) for attention and frequent experts with uniform-bits quantization (i.e. $3$ and $8$ bits). The Pareto-optimal solution is $3.29$ bits.
4.3. Benchmark Results
We first evaluate the four bit-allocation heuristic MoE quantization approaches using GPTQ on Mixtral-8x7B and DeepSeek-MoE-16B.
Table 1. Comparison of quantization bit allocation strategies including attention (Attn) and shared expert (Shared) prioritization, frequency-based expert selection (Freq), first-layer prioritization (FirstL), linear outlier score predictor (LinearOSP) and layer importance score predictor (LayerISP). We evaluate performance on Mixtral-8x7B and DeepSeek-MoE-16B-base across multiple benchmarks. Unlike uniform-precision quantization (i.e. GPTQ), our approach tailors bit allocation to the MoE structure, demonstrating the effectiveness of structure-aware precision allocation, especially in low-bit settings.
Mixtral-8x7B
GPTQ $3$ $73.56$ $93.00$ $44.80$ $75.31$ $79.11$ $58.37$ $70.69$
GPTQ $2$ $49.33$ $63.00$ $25.40$ $28.18$ $52.99$ $24.29$ $40.53$
+Attn $2.06$ $54.14$ $65.00$ $26.80$ $44.87$ $57.78$ $27.38$ $46.00$5.47
+Attn+Freq $3.27$ $73.09$ $92.00$ $44.80$ $77.03$ $78.29$ $63.05$ $71.38$30.85
+Attn+Freq+FirstL $3.51$ $73.32$ $95.00$ $44.00$ $80.06$ $79.71$ $64.89$ $72.83$32.30
+Attn+LinearOSP $3.12$ $72.85$ $92.00$ $43.40$ $78.14$ $78.40$ $62.92$ $71.29$30.76
+Attn+LayerISP $3.35$ $71.19$ $93.00$ $44.00$ $77.03$ $78.29$ $63.15$ $71.11$30.58
DeepSeek-MoE-16B-base
GPTQ $3$ $67.87$ $87.00$ $40.40$ $72.92$ $78.50$ $39.12$ $64.30$
GPTQ $2$ $53.27$ $76.00$ $30.20$ $45.32$ $66.53$ $25.28$ $49.43$
+Attn $2.06$ $61.72$ $83.00$ $34.00$ $61.93$ $73.45$ $26.53$ $56.77$7.34
+Attn+Shared $2.12$ $65.67$ $85.00$ $39.20$ $67.84$ $76.28$ $35.28$ $61.55$12.12
+Attn+Shared+Freq $3.00$ $68.82$ $88.00$ $42.00$ $73.32$ $77.31$ $41.51$ $65.16$15.90
+Attn+Shared+Freq+FirstL $3.06$ $68.35$ $88.00$ $42.60$ $73.85$ $77.69$ $41.58$ $65.35$15.92
+Attn+Shared+LinearOSP $3.06$ $67.32$ $88.00$ $41.60$ $72.45$ $77.48$ $41.64$ $64.84$15.41
+Attn+Shared+LayerISP $3.06$ $69.30$ $87.00$ $42.00$ $73.56$ $77.86$ $41.86$ $65.27$15.67
Table 1 presents the overall comparison performance of our mixed-precision MoE quantization strategies v.s. baseline (i.e. GPTQ) across multiple benchmarks Mixtral-8x7B and DeepSeek-MoE-16B-base. The results highlight that uniform quantization leads to substantial performance degradation, reinforcing the need for structure-aware, fine-grained quantization strategies. By incorporating methods such as attention-aware adjustments (+Attn), expert frequency-based bit allocation (+Freq), and prioritization of early MoE layers (+FirstL), importance-based quantization (+LinearOSP and +LayerISP), we observe improvements in model performance, particularly in lower-bit settings. A key takeaway from the results is that a mixed-precision approach tailored to the MoE structure outperforms fixed-bit quantization. In particular, expert usage frequency-based quantization (+Freq) improves performance by dynamically allocating more bits to frequently used experts, though its impact varies depending on the routing balance of different models. Additionally, prioritizing the first few MoE blocks (+FirstL) yields better results. Lastly, importance-based strategies(+LinearOSP, +LayerISP) further refine bit allocation, demonstrating the effectiveness of adaptive quantization at different levels of granularity.
To gain deeper insights, we further systematically analyze four key research questions: ($1$) the effectiveness of expert usage frequency as a quantization heuristic, ($2$) whether attention or FFNN layers deserve more bit precision, ($3$) the importance of the first versus last MoE blocks in quantization, and ($4$) the necessity of allocating more bits to shared experts. The following sections provide a detailed discussion of each aspect.